# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ 7 Concluding Remarks
In this paper, we present a systematic study of how LRS influences scaling laws by considering a teacher–student kernel regression setting. Our theoretical framework yields a novel functional scaling law, which explicitly characterizes the impact of both learning rate and batch size schedules via a convolution-type functional term. The utility of our FSL is demonstrated through detailed analyses of three widely used LRSs, providing theoretical justification for several prevailing practices in LLM pre-training – most notably, offering an explanation for the effectiveness of the empirically popular but previously less-understood WSD schedules.
Looking ahead, several promising directions remain open. First, it is essential to develop a theoretical characterization of the optimal LRS under various resource constraints, which remains largely unresolved. Second, leveraging FSL studies the effect of batch size scheduling – a relatively under-explored topic despite its practical importance. Third, to make FSL truly practical, it is crucial to refine our FSL through extensive large-scale LLM pre-training experiments.
@@ Acknowledgement
Lei Wu is supported by the National Natural Science Foundation of China (NSFC12522120, NSFC92470122, NSFC12288101). Binghui Li is supported by the Elite Ph.D. Program in Applied Mathematics for PhD Candidates in Peking University. We would like to thank Kaifeng Lyu, Kairong Luo, Haodong Wen, Tingkai Yan, Yuhao Liu, Yunze Wu and Zean Xu for helpful discussions and anonymous reviewers for their valuable suggestions.
@@
Appendix A Additional Related Work
A.1 Empirical Findings on Scaling Laws
Scaling laws have emerged as a foundational principle in understanding the interplay between model size, dataset size, and performance in large-scale deep model pre-training. Initially proposed by [29], these laws have been further refined through a series of subsequent studies [22, 24, 70, 28, 1, 9, 57, 31, 60, 42].
Among these, [24] introduced the Chinchilla Law, which approximates the final-step risk $\mathcal{R}$ as follows:
$\mathcal{R}=R_{0}+A_{1}D^{-\kappa_{1}}+A_{2}M^{-\kappa_{2}},$
where $R_{0}$, $A_{1}$, $A_{2}$, $\kappa_{1}$, and $\kappa_{2}$ are constants, and $D$ and $M$ represent the amount of training data (tokens) and model size (number of parameters), respectively. Based on this risk prediction, they formulate the following compute-optimal scaling problem:
$\operatornamewithlimits{min}_{DM=C}R_{0}+A_{1}D^{-\kappa_{1}}+A_{2}M^{-\kappa_{2}},$
where the compute budget $C=DM$ serves as the scaling bottleneck.
Furthermore, [60] and [42] empirically investigated scaling laws in the context of LRSs. [60] heuristically proposed the Momentum Law to characterize the impact of LRSs on scaling laws, where the population risk at the $k$-th step follows:
$\mathcal{R}_{k}=R_{0}+AS_{1}^{-\kappa}-CS_{2},$
where $R_{0}$, $A$, $C$, and $\kappa$ are constant parameters, and $S_{1}$ and $S_{2}$ roughly denote the area under the learning rate curve and the learning rate annealing area, respectively, which are formally defined as follows:
$S_{1}:=\sum_{i=1}^{k}\eta_{i},\quad\quad S_{2}:=\sum_{i=2}^{k}\sum_{j=2}^{i}(\eta_{j-1}-\eta_{j})\lambda^{i-j},$
where $\bm{\eta}:=(\eta_{1},\eta_{2},\cdots,\eta_{K})^{\top}\in\mathbb{R}_{+}^{K}$ denotes the learning rate schedule and $\lambda$ is a hyper-parameter representing the decay factor for LR annealing momentum, which typically ranges from $0.99$ to $0.999$.
Subsequently, [42] further developed an empirical law, termed the Multi-Power Law, in which they replaced $S_{2}$ in the Momentum Law with additional power laws to account for the loss reduction effect as the learning rate decays. The Multi-Power Law says that the risk can be predicted as:
$\mathcal{R}_{k}=R_{0}+AS_{1}^{-\kappa}-\operatorname{LD}(k),$
$LD(k)=C\sum_{i=2}^{k}(\eta_{i-1}-\eta_{i})$ $G(\eta_{i}^{-\kappa^{\prime}}S_{i}),\quad\quad S_{i}=\sum_{j=1}^{i}\eta_{j},\quad\quad G(x)=1-(C^{\prime}x+1)^{-\kappa^{\prime\prime}},$
where $R_{0},A,C,C^{\prime},\kappa,\kappa^{\prime},\kappa^{\prime\prime}$ are constants.
However, both the Momentum Law and the Multi-Power Law remain empirical, lacking a rigorous theoretical foundation. In our work, we build on a solvable model and theoretically derive a novel functional-form scaling law with arbitrary LRS, which also generalizes well for unseen larger model sizes.
A.2 SGD for Kernel Regression
Several works have analyzed the SGD algorithm within the context of high-dimensional kernel regression [16, 38, 53, 47, 8, 62]. [16] demonstrated that averaged-iteration SGD with a constant learning rate can achieve statistical optimality in the easy-task regime. [38] showed that multiple-pass, non-averaged SGD also attains statistical optimality in this regime, while [53] established that multiple-pass, averaged-iteration SGD can achieve optimality in the hard-task regime. Building on these results, [47] analyzed the convergence behavior of SGD in high-dimensional kernel regression, demonstrating that strategies such as averaging and multiple passes can enhance learning performance. [8] and [62] studied the noiseless case: [8] analyzed SGD for a noiseless linear model, deriving convergence results and extending the findings to online function estimation. [62] examined SGD in a noiseless least-squares setup, showing convergence for non-strongly convex problems and faster-than-usual convergence rates in over-parameterized settings.
Appendix B Experiment Details and Additional Results
B.1 Popular Learning Rate Schedules
@@
B.2 Teacher-Student Kernel Regression
@@
In each experiment, we adopt a teacher-student model configuration with $M=N=128$, $\sigma=3$ and employ the top-$M$ projection matrix, thereby eliminating the approximation error term $M^{-s\beta}$. We explore a range of values for $s\in[0.5,1]$ and $\beta\in[1.5,5]$, encompassing both easy- ($s\geqslant 1-1/\beta$) and hard-learning ($s<1-1/\beta$) regimes. For each parameter configuration, we execute 200 independent SGD runs with a batch size of 1 over 10,000 steps. The resulting average trajectory across these runs serves as the fitting target. The FSL fitting is performed using the physical-time FSL formulation.
$\mathcal{E}_{k}=c_{1}T(k)^{-s}+c_{2}\int_{0}^{k}\mathcal{K}(T(k)-T(u))e(T(u))\varphi(u)^{2}\mathop{}\!\mathrm{d}u+c_{3}\sigma^{2}\int_{0}^{k}\mathcal{K}(T(k)-T(u))\varphi(u)^{2}\mathop{}\!\mathrm{d}u,$
where $c_{1}$, $c_{2}$, $c_{3}$ are constants to fit.
Fitting the SGD trajectory, we minimize the mean squared error (MSE) between the empirical risk trajectory of SGD, denoted by $\mathcal{E}_{\text{SGD}}(k)$, and the theoretical prediction from FSL, $\mathcal{E}_{k}$. Formally, we solve the following optimization problem:
@@
  • FSL experiments. This experiment is intended to validate the theoretical predictions derived from the FSL. We compute the predicted risk by numerically discretizing the FSL (7), with all untracked constants set to 1. For each LRS, following the theoretical analysis, we set $\eta_{\max}=0.05D^{-r}$, where $r=s/(1+s)$ for the constant learning rate schedule, and $r=1$ for exponential decay and WSD schedules. We fix the batch size to $B=1$; thus, for each data budget $D$, we compute the intrinsic time and evaluate the final-step risk using the discretized FSL.
@@ B.3 LLM pre-training
In this section, in order to fit the scaling laws on real LLM pre-training loss curves, we will derive an approximation form of the FSL in Theorem 4.1.
@@
$\int_{T(i)}^{T(k)}\mathcal{K}(T(k)-t)\mathop{}\!\mathrm{d}t=\int_{M^{-\beta}}^{1}u^{1-1/\beta}\int_{T(i)}^{T(k)}e^{-2u(T(k)-t)}\mathop{}\!\mathrm{d}t\mathop{}\!\mathrm{d}u=\frac{1}{2}-\frac{1}{2}g_{1-1/\beta}(T(k)-T(i)),$
we then further simplify it as (see the definition of $g$ function in Section C.3)
$\mathrm{LRD}(k)$ $\approx\frac{1}{2B}\sum_{i=1}^{k}(\eta_{i-1}-\eta_{i})(\sigma^{2}+e(\xi_{i}))(1-g_{1-1/\beta}(T(k)-T(i)))$
$\approx\frac{1}{2B}\sum_{i=1}^{k}(\eta_{i-1}-\eta_{i})(\sigma^{2}+e(T(i)))(1-(1+c(T(k)-T(i)))^{-\gamma}).$
Here, in the last line we approximate $\xi_{i}$ as $T(i)$ and approximate $g_{1-1/\beta}(t)$ as $(1+ct)^{-\gamma}$ with some constant $c>0$ for simplicity.
@@
The objective of this experiment is to analyze and fit the loss function using our functional scaling law, by (12),
$\mathcal{L}_{\Theta}(k)=L_{0}+\frac{c_{1}}{T(k)^{s}}+c_{2}M^{-s\beta}-\operatorname{LRD}(k)$
where
$\operatorname{LRD}(k):=c_{3}\sum_{i=1}^{k}(\eta_{i-1}-\eta_{i})\left(c_{4}+\frac{1}{T(i)^{s}}\right)\left(1-\frac{1}{(1+c_{5}(T(k)-T(i)))^{\gamma}}\right),$
and $\Theta=(L_{0},c_{1},c_{2},c_{3},c_{4},c_{5},s,\beta,\gamma)$.
@@
We propose to obtain a numerical optimal LRS by directly minimizing the final-step risk over the space of LRS using the fitted FSL, termed FSL-optimal LRS.
@@
We further validate the applicability of our functional scaling law (FSL) across varying training durations. Using a 100M LLaMA dense model, we conduct experiments with total training steps set to 17k, 34k, 68k, and 134k. As demonstrated in Figures 7 and 8, our FSL accurately models the loss trajectories across all evaluated step counts, confirming its robustness to different total training steps.
Finally, we conduct a comprehensive empirical comparison between our FSL-optimal learning rate schedule and various WSD baselines, examining different decay ratios and minimum learning rate configurations. As evidenced by Figures 9 and 10, our FSL-optimal LRS consistently outperforms all WSD variants, achieving superior final pre-training loss across all experimental conditions. This systematic evaluation demonstrates both the effectiveness of our theoretically-derived schedule and its practical advantages over conventional heuristic approaches.

### v3

@@ 7 Concluding Remarks
In this paper, we present a systematic study of how LRS influences scaling laws by considering a power-law kernel regression problem. Our theoretical framework yields a novel functional scaling law, which explicitly characterizes the impact of both learning rate and batch size schedules via a convolution-type functional term. The utility of our FSL is demonstrated through detailed analyses of three widely used LRSs, providing theoretical justification for several prevailing practices in LLM pre-training—most notably, offering an explanation for the effectiveness of the empirically popular but previously less-understood WSD schedules.
Looking ahead, several promising directions remain open. First, it is essential to develop a theoretical characterization of the optimal LRS under various resource constraints, which remains largely unresolved. Second, leveraging FSL studies the effect of batch size scheduling—a relatively under-explored topic despite its practical importance. Third, to make FSL truly practical, it is crucial to refine our FSL through extensive large-scale LLM pre-training experiments.
@@ Acknowledgement
Lei Wu is supported by the National Natural Science Foundation of China (NSFC12522120, NSFC92470122, and NSFC12288101). Binghui Li is supported by the Elite Ph.D. Program in Applied Mathematics at Peking University. We are grateful to Kaifeng Lyu, Kairong Luo, and Haodong Wen for generously sharing their work [40], which greatly inspired this study. We also thank Tingkai Yan, Yuhao Liu, Yunze Wu, and Zean Xu for many helpful discussions, and the anonymous reviewers for their valuable feedback.
@@
Appendix A Miscellanea
A.1 Empirical Fitting of LLM Pre-training Loss Trajectory
The Chinchilla Law [21] describes the final-step loss $\mathcal{L}$ as follows:
$L(M,D)=L_{0}+A_{1}M^{-\kappa_{1}}+A_{2}D^{-\kappa_{2}},$
where $L_{0}$, $A_{1}$, $A_{2}$, $\kappa_{1}$, and $\kappa_{2}$ are constants, and $D$ and $M$ represent the amount of training data (tokens) and model size (number of parameters), respectively.
Later, [60] proposed the Momentum Law, a heuristic rule designed to capture the full loss trajectory. Given a learning rate schedule $\bm{\eta}:=\{\eta_{j}\}_{j}$, the loss at the $k$-th step is modeled as
$\mathcal{L}_{k}(\bm{\eta})$ $=L_{0}+AS_{1}^{-\kappa}-CS_{2},$
where
$S_{1}=\sum_{i=1}^{k}\eta_{i},\quad S_{2}=\sum_{i=2}^{k}\sum_{j=2}^{i}(\eta_{j-1}-\eta_{j})\lambda^{i-j}.$
Here $L_{0}$, $A$, $C$, and $\kappa$ are constants, and $\lambda\in(0,1)$ is a hyperparameter representing the decay factor for learning rate annealing, which typically ranges from $0.99$ to $0.999$.
Subsequently, [40] proposed the Multi-Power Law (MPL), which replaces the $S_{2}$ in the Momentum Law with additional power laws to better capture the progressive loss reduction induced by learning-rate decay. Specifically, the MPL takes the following form:
$\mathcal{L}_{k}(\bm{\eta})=L_{0}+AS_{1}^{-\kappa}-\operatorname{LD}(k),$ (27)
where
$\operatorname{LD}(k):=C\sum_{i=2}^{k}(\eta_{i-1}-\eta_{i})G(\eta_{i}^{-\kappa^{\prime}}S_{i}),\quad S_{i}:=\sum_{j=1}^{i}\eta_{j},\quad G(x):=1-(C^{\prime}x+1)^{-\kappa^{\prime\prime}}.$
Here $L_{0},A,C,C^{\prime},\kappa,\kappa^{\prime},\kappa^{\prime\prime}$ are constants.
A.2 Popular Learning Rate Schedules in LLM Pre-training
@@
A.3 Connections to Kernel Regression
In this section, we explain how our setup in Section 2 are equivalent to learning with kernels.
Definition A.1 (Positive semidefinite (PSD) kernel).
A function $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is called a continuous positive semidefinite (PSD) kernel if it satisfies:
  • Symmetry: $K(\mathbf{x},\mathbf{x}^{\prime})=K(\mathbf{x}^{\prime},\mathbf{x})$ for all $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}$;
  • Positive semidefiniteness: for any $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\mathcal{X}$ and $a_{1},\dots,a_{n}\in\mathbb{R}$,
    $\sum_{i,j=1}^{n}a_{i}a_{j}\,K(\mathbf{x}_{i},\mathbf{x}_{j})\geqslant 0.$
Definition A.2 (Reproducing kernel Hilbert space (RKHS)).
Given a kernel $K$, the reproducing kernel Hilbert space $\mathcal{H}_{K}$ associated with $K$ is a Hilbert space of functions $f:\mathcal{X}\to\mathbb{R}$ such that
$\langle f,K(\mathbf{x},\cdot)\rangle_{\mathcal{H}_{K}}=f(\mathbf{x}),\quad\forall f\in\mathcal{H}_{K},~\mathbf{x}\in\mathcal{X}.$
Kernel methods learn functions from a hypothesis space defined by the associated RKHS. For instance, kernel ridge regression gives estimator:
$\widehat{f}_{\lambda}=\operatorname*{arg\,min}_{f\in\mathcal{H}_{K}}\frac{1}{n}\sum_{i=1}^{n}(f(\mathbf{x}_{i})-y_{i})^{2}+\lambda\|f\|_{\mathcal{H}_{K}}^{2}.$
Hence, model capacity is determined by the size of the RKHS $\mathcal{H}_{K}$.
Let $\mathcal{D}$ be the input distribution. Given a kernel $K$, define the associated integral operator $\mathcal{T}_{K}:L^{2}(\mathcal{D})\to L^{2}(\mathcal{D})$ by
$\mathcal{T}_{K}f(\cdot)=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[K(\cdot,\mathbf{x})f(\mathbf{x})].$
By assuming $\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[K(\mathbf{x},\mathbf{x})]<\infty$, the operator $\mathcal{T}_{K}$ is compact (Mercer’s theorem) and consequently, the kernel admits the following eigenvalue decomposition
$K(\mathbf{x},\mathbf{x}^{\prime})=\sum_{j=1}^{\infty}\lambda_{j}e_{j}(\mathbf{x})e_{j}(\mathbf{x}^{\prime}),$
where $\{\lambda_{j}\}_{j=1}^{\infty}$ and $\{e_{j}\}_{j=1}^{\infty}$ denotes the eigenvalues and eigenfunctions, respectively. Moreover, $\langle e_{i},e_{j}\rangle_{L^{2}(\mathcal{D})}=\delta_{i,j}$, i.e., the eigenfunctions form an orthonormal basis of $L^{2}(\mathcal{D})$.
Using the spectral decomposition, the RKHS admits the following representation:
$\mathcal{H}_{K}=\left\{\sum_{j=1}^{\infty}a_{j}e_{j}:\sum_{j=1}^{\infty}\frac{a_{j}^{2}}{\lambda_{j}}<\infty\right\}.$
To better quantify the smoothness of functions, we often consider the interpolation space $\mathcal{H}_{K}^{s}$ with $s\geqslant 0$, defined as
$\mathcal{H}_{K}^{s}=\left\{\sum_{j=1}^{\infty}a_{j}e_{j}:\sum_{j=1}^{\infty}\frac{a_{j}^{2}}{\lambda_{j}^{s}}<\infty\right\}.$
Clearly, $\mathcal{H}_{K}^{1}=\mathcal{H}_{K}$, and
$\mathcal{H}_{K}^{s_{1}}\subset\mathcal{H}_{K}^{s_{2}},\qquad\forall\,s_{1}>s_{2}\geqslant 0.$
Hence, the index $s$ characterizes the smoothness of a function relative to the chosen kernel.
In the analysis of kernel methods, the following conditions are commonly used to describe the smoothness of the target function and the capacity of the kernel, respectively.
Assumption A.3 (Source condition).
There exists some $s>0$ such that $f^{*}\in\mathcal{H}_{K}^{s}$.
Assumption A.4 (Capacity condition).
There exists some $\beta>1$ such that $\lambda_{j}\eqsim j^{-\beta}$.
These conditions yield the following interpretation:
  • A smaller $s$ indicates that the target function $f^{*}$ belongs to a larger space, corresponding to a more difficult learning problem.
  • A smaller $\beta$ implies a slower eigenvalue decay, meaning a richer hypothesis space $\mathcal{H}_{K}$ and thus higher model capacity.
Our formulation in Section 2 is equivalent to the above setting, but expressed in terms of the feature map $\bm{\phi}$. Under Assumption 2.2, we have
$K_{\phi}(\mathbf{x},\mathbf{x}^{\prime})=\sum_{j=1}^{N}\phi_{j}(\mathbf{x})\phi_{j}(\mathbf{x}^{\prime})=\sum_{j=1}^{N}\lambda_{j}\,\widehat{\phi}_{j}(\mathbf{x})\widehat{\phi}_{j}(\mathbf{x}^{\prime}).$
In this case, Assumption 2.3 corresponds exactly to the above capacity condition, while the task-difficulty assumption in Section 2 can be viewed as a power-law version of the source condition. Specifically, under Assumption 2.5,
$f^{*}=\sum_{j=1}^{N}\theta_{j}^{*}\phi_{j}=\sum_{j=1}^{N}j^{-1/2}\lambda_{j}^{\,s}\,\widehat{\phi}_{j}=\sum_{j=1}^{N}a_{j}\,\widehat{\phi}_{j}.$
Hence, for any arbitrarily small $\delta\in(0,1)$, we have $f^{*}\in\mathcal{H}_{K_{\phi}}^{s-\delta}$, since
$\sum_{j=1}^{N}\frac{a_{j}^{2}}{\lambda_{j}^{s-\delta}}=\sum_{j=1}^{N}j^{-1-\beta(s-\delta)}<\infty.$
A.4 The SDE Modeling
The physical-time SDE.
In our setup, the SGD update can be written as
$\mathbf{v}_{k+1}=\mathbf{v}_{k}-\varphi_{k}\nabla\mathcal{R}(\mathbf{v}_{k})h-\varphi_{k}h\bm{\xi}_{k}.$
The term $\bm{\xi}_{k}$ is the gradient noise, whose covariance is $\frac{1}{B_{k}}\Sigma(\mathbf{v}_{k})$. By assuming the gradient noise to be Gaussian, the SGD becomes
$\mathbf{v}_{k+1}-\mathbf{v}_{k}=-\varphi_{k}\nabla\mathcal{R}(\mathbf{v}_{k})h+\varphi_{k}\sqrt{h}\,\,\mathcal{N}\left(0,\tfrac{h}{B_{k}}\Sigma(\mathbf{v}_{k})\right).$
It is exactly the Euler–Maruyama discretization of the Itô-type SDE:
where $\mathbf{B}_{\tau}\in\mathbb{R}^{M}$ denotes the $M$-dimensional Brownian motion, and $\varphi(\cdot)$, $b(\cdot)$ are the continuous version of LRS function and batch-size schedule function, respectively.
The intrinsic-time SDE.
Intuitively, the discrete update (5) can be viewed as the Euler–Maruyama discretization of SDE (8) on the non-uniform grid $\{t_{k}=\sum_{j=0}^{k}\eta_{j}\}_{k\in\mathbb{N}}$ where the effective step size is $\Delta t_{k}=\eta_{k}$:
$\mathbf{v}_{k+1}-\mathbf{v}_{k}=-\nabla\mathcal{R}(\mathbf{v}_{k})(t_{k+1}-t_{k})-\sqrt{t_{k+1}-t_{k}}\;\mathcal{N}\left(0,\tfrac{\eta_{k}}{B_{k}}\Sigma(\mathbf{v}_{k})\right).$
Appendix B Experiment Details and Additional Results
B.1 Power-Law Kernel Regression
@@
In each experiment, we adopt a PLKR configuration with $M=N=128$, $\sigma=3$ and employ the top-$M$ projection matrix, thereby eliminating the approximation error term $M^{-s\beta}$. We explore a range of values for $s\in[0.5,1]$ and $\beta\in[1.5,5]$, encompassing both easy- ($s\geqslant 1-1/\beta$) and hard-learning ($s<1-1/\beta$) regimes. For each parameter configuration, we execute 200 independent SGD runs with a batch size of 1 over 10,000 steps. The resulting average trajectory across these runs serves as the fitting target. The FSL fitting is performed using the physical-time FSL formulation.
$\mathcal{E}_{k}=c_{1}T(k)^{-s}+c_{2}\sum_{i=1}^{k}\mathcal{K}(T(k)-T(i))e(T(i))\eta_{i}^{2}+c_{3}\sigma^{2}\sum_{i=1}^{k}\mathcal{K}(T(k)-T(i))\eta_{i}^{2},$
where $c_{1}$, $c_{2}$, $c_{3}$ are constants to fit, $\{\eta_{i}\}_{i=1}^{k}$ is the learning rate schedule, and $T(i)=\sum_{j=1}^{i}\eta_{j}$.
When fitting the SGD trajectory, we minimize the mean squared error (MSE) between the empirical risk trajectory of SGD (without the irreducible risk $\frac{\sigma^{2}}{2}$), denoted by $\mathcal{E}_{\text{SGD}}(k)$, and the theoretical prediction from FSL, $\mathcal{E}_{k}$. Formally, we solve the following optimization problem:
@@
  • FSL experiments. This experiment is intended to validate the theoretical predictions derived from the FSL. We compute the predicted risk by numerically discretizing the FSL (11), with all untracked constants set to 1. For each LRS, following the theoretical analysis, we set $\eta_{\max}=0.05D^{-r}$, where $r=s/(1+s)$ for the constant learning rate schedule, and $r=1$ for exponential decay and WSD schedules. We fix the batch size to $B=1$; thus, for each data budget $D$, we compute the intrinsic time and evaluate the final-step loss using the discretized FSL.
@@ B.2 LLM pre-training
In this section, in order to fit real LLM pre-training loss curves, we will derive an approximation form of the FSL in Theorem 4.2.
@@
$\int_{T(i)}^{T(k)}\mathcal{K}(T(k)-t)\mathop{}\!\mathrm{d}t\approx\int_{T(i)}^{T(k)}\frac{1}{(1+ct)^{2-1/\beta}}\mathop{}\!\mathrm{d}t\mathop{}\!\mathrm{d}u\eqsim 1-\frac{1}{(1+c(T(k)-T(i))^{1-1/\beta}},$
we then further simplify it as
$\mathrm{LRD}(k)$ $\approx\frac{1}{B}\sum_{i=1}^{k}(\eta_{i-1}-\eta_{i})(\sigma^{2}+e(T(i)))(1-(1+c(T(k)-T(i)))^{-\gamma}).$
Here, we approximate $\xi_{i}$ as $T(i)$ and introduce a new parameter $\gamma$ to replace $1-\frac{1}{\beta}$ for simplicity.
@@
The objective of this experiment is to analyze and fit the loss function using our functional scaling law, by (30), since we do not explore the effect of varying the model size $M$ in our experiments, we drop the term $M^{-s\beta}$ and get
$\mathcal{L}_{\Theta}(k)=L_{0}+\frac{c_{1}}{T(k)^{s}}-\operatorname{LRD}(k)$
where $T(k)=\sum_{i=1}^{k}\eta_{i}$ and
$\operatorname{LRD}(k):=c_{2}\sum_{i=1}^{k}(\eta_{i-1}-\eta_{i})\left(c_{3}+\frac{1}{T(i)^{s}}\right)\left(1-\frac{1}{(1+c_{4}(T(k)-T(i)))^{\gamma}}\right),$
and $\Theta=(L_{0},c_{1},c_{2},c_{3},c_{4},s,\gamma)$.
@@
We propose to obtain a numerical optimal LRS by directly minimizing the final-step loss over the space of LRS using the fitted FSL, termed FSL-optimal LRS.
@@
We further validate the applicability of our functional scaling law (FSL) across varying training durations. Using a 100M LLaMA dense model, we conduct experiments with total training steps set to 17k, 34k, 68k, and 134k. As demonstrated in Figures 6 and 7, our FSL accurately models the loss trajectories across all evaluated step counts, confirming its robustness to different total training steps.
Finally, we conduct a comprehensive empirical comparison between our FSL-optimal learning rate schedule and various WSD baselines, examining different decay ratios and minimum learning rate configurations. As evidenced by Figures 8 and 9, our FSL-optimal LRS consistently outperforms all WSD variants, achieving superior final pre-training loss across all experimental conditions. This systematic evaluation demonstrates both the effectiveness of our theoretically-derived schedule and its practical advantages over conventional heuristic approaches.